# Optimizing Multi-Modal Trackers via Sensitivity-aware Regularized Tuning

### v1

@@ IV-C Ablation Study and Analysis
Effectiveness of Proposed Sensitivities. We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed sensitivity items, as summarized in Tab. IV. A comparison of (a) and (b) demonstrates that fine-tuning significantly enhances the domain adaptation ability. Further, the contrast between (b) and (c) (or (d) and (e)) figures out that the prior sensitivity regularization scheme significantly improves SR by $2.4\%$ and PR by $3.2\%$ on LasHeR, highlighting its effectiveness. Similarly, the comparison between (b) and (d) (or (c) and (e)) reveals that rebalancing gradients optimized with diverse transfer data leads to substantial improvements. Notably, the simultaneous application of both prior and transfer sensitivity yields significantly greater improvements than either method alone, proving their complementary nature. While the transfer sensitivity penalty alone yields only modest improvement on FE108, it still contributes to the overall leading performance of our method.
TABLE IV: Ablative study results of the proposed key components. “FFT” refers to full fine-tuning of the backbone; (a) is the zero-shot performance; (b) serves as our baselines; (e) denotes the complete regularized tuning framework.
Exp. FFT Prior Sensitivity Transfer Sensitivity FE108 DepthTrack LasHeR
@@
Eigenvalue Budget (Top-$K$) for Prior Sensitivity. In this section, we examine the effect of the low-dimension eigenvalue count $K$ in Eq. (8). For each operation, we record up to the 20 largest eigenvalues during Rayleigh-quotient probing and compute the prior sensitivity using the top-$K$ of them. As shown in Tab. V, decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best configuration), and even $K=1$ still yields a clear improvement over the baseline. This indicates that the effectiveness of our regularizer depends primarily on the relative ranking of sensitivities rather than their absolute magnitudes. Under our low-dimension approximation, the central challenge lies in accurately identifying the operations most sensitive to the generalization gap.
@@
Settings of Sensitivity Harmony Coefficient $\kappa$. In this section, we evaluate the impact of different sensitivity harmony coefficients, $\kappa$ as defined in Eq. (14), with the detailed results presented in Tab. VI. By systematically increasing $\kappa$ from 0.1 to 0.9, we observe that a moderately large coefficient (e.g., 0.6) yields the best performance, outperforming both smaller (e.g., 0.3) and larger values (e.g., 0.9). This finding suggests that while both prior sensitivity and transfer sensitivity contribute to performance improvements from different perspectives, a delicate balance between them further optimizes the fine-tuning process, promoting better generalization and adaptability.
@@
Impact of Box Head Tuning. In our default setting, “full fine-tuning” updates the backbone while keeping the box head frozen. To assess the impact of tuning the box head, we also examine a more aggressive variant, complete fine-tuning, where the box head is updated as well. As shown in Tab. VII, unfreezing the box head further exacerbates over-fitting. While our regularization method remains effective, the results reveal an important caveat: unfreezing the box head during cross-modal adaptation can be risky, as it tends to disrupt the modal-agnostic object association knowledge encoded during pre-training.
@@
Compatibility with PEFT Methods. Existing PEFT methods [15, 16, 17] typically freeze pre-trained parameters and update only a minimal number of additional parameters, which can hinder sufficient optimization. To evaluate the compatibility of our proposed regularization techniques with these methods, we unfreeze their backbones and retrain them with our regularization applied. As shown in Tab. VIII, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR. Notably, even for the unified tracker UnTrack, our approach achieves superior results on LasHeR: $2.7\%$ in SR and $3.7\%$ in PR. This demonstrates that overly rigid constraints on pre-trained models limit their transfer potential. However, our method negatively impacts SDSTrack, likely because it optimizes the pre-trained parameters, whereas SDSTrack relies on modal-specific adapters trained from scratch.
@@
Compatibility with Single-modality Methods. This work aims to mitigate the misfitting issue when adapting the foundation trackers to downstream tasks. A central question is how the proposed regularization methods perform on single-modality data. To investigate this, we conduct ablation studies highlighting their impact on different modalities. As shown in Tab. IX, both RGB and auxiliary modalities benefit significantly from our regularization techniques. For example, the RGB and depth modalities achieve F-score gains of $5.3\%$ and $4.0\%$, respectively, on the DepthTrack dataset. Despite substantial distribution differences, our method significantly and consistently enhances the adaptability of auxiliary modalities across various tasks. These findings underscore the importance and necessity of applying constraints when transferring the pre-trained trackers to downstream domains.
@@
Impact of Event Representations. In this work, we focus on constructing suitable frame-like event representations tailored to cross-modal transfer. In this experiment, we keep the same training setup, varying only the input event representations. As shown in Tab. X, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps. In contrast, count-based representations (Event Count, Event Frame) offer more robust performance. The time-interpolation-based Event Volume further improves results by utilizing time-weighted and multi-channel methods to preserve spatio-temporal information. Notably, we employ simple color-polarity event images (with ViPT) to align with the RGB pre-trained model, leading to superior performance.
@@
Effectiveness of Smaller Learning Rates. In this paper, we exploit the sensitivity-regularized gradients to bolster the transfer process. A straightforward strategy might involve using a smaller learning rate. To investigate this, we evaluate the effect of reduced learning rates. As shown in Tab. XI, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR. Further reduction (i.e., $\alpha=10^{-6}$) leads to overall performance degradation. These findings suggest that merely reducing the learning rate is insufficient to address over-fitting in cross-modality adaptation, while simultaneously limiting transfer potential, as it uniformly suppresses parameter updates without prioritizing those that are sensitive or high-risk.
@@
Exp. SR PR
$\alpha=10^{-4}$ 53.2 65.8
$\alpha=10^{-5}$ 52.9 66.2
$\alpha=10^{-6}$ 52.5 65.4
Ours 56.3 70.1
@@
TABLE XII: Computational complexity and speed analysis on the LasHeR dataset.
@@ IV-D Computational Efficiency
Efficiency of Regularized Fine-tuning. We compare the training efficiency of our method against vanilla fine-tuning. Since prior sensitivity estimation is conducted as a one-time preprocessing step before fine-tuning, its runtime is excluded from the training time. Our method introduces parameter-wise transfer sensitivity quantification with marginal computational overhead, as this sensitivity is derived from the off-the-shelf gradients. While vanilla fine-tuning operates at 37.5 ms per iteration, our method runs at 39.0 ms per iteration, incurring only a modest $4.0\%$ increase in training time.
Inference Cost and Speed. We compare the computational complexity and runtime efficiency of existing methods, focusing on models built upon OSTrack, DropTrack and SUTrack. Note that the proposed regularization techniques are applied only during training, imposing no additional computational burden during testing. All experiments are conducted on the same system, which includes an Intel(R) Xeon(R) 6456C CPU, 256 GB RAM, and one NVIDIA 3090Ti GPU. As reported in Tab. XII, our method supports real-time tracking at 24.8 frames per second, while delivering top-tier performance. When integrated into ViPT, our approach (“ViPT+Ours”) yields both superior speed and performance. Under the DropTrack base model, it still operates at 48.8 frames per second, maintaining a respectable speed given its focus on accuracy.
IV-E Visual Comparison
@@ V Conclusion
Proof of Proposition 1.
  1. 1.
    Bounded FIM Error. Each group approximation $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}$ captures the principal tangent of $\mathcal{F}^{(\theta_{0}^{j})}$ with bounded error. Due to symmetry and orthogonal eigenvectors, the Frobenius norm of the error is derived from the difference in their eigenvalues:
@@
  2. 2.
    Bounded Generalization Gap Error. The approximate FIM $\tilde{\mathcal{F}}^{(\theta_{0})}$ remains close to $\mathcal{F}^{(\theta_{0})}$ as a metric on parameter space. For any parameter difference $\Delta\theta\in\mathbb{R}^{|\theta|}$, let $\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\mathcal{F}^{(\theta_{0})}\Delta\theta$ denote the generalization gap induced by the true FIM. Likewise $\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\tilde{\mathcal{F}}^{(\theta_{0})}\Delta\theta$ is the approximated generalization gap. Then the discrepancy between these distances is bounded in terms of the residual generalization gap error. In particular, one has following formula by definition of the spectral norm:

### v2

@@ IV-C Ablation Study and Analysis
Effectiveness of Proposed Sensitivities. We conduct a series of comprehensive experiments to uncover the interplay and effectiveness of the two proposed sensitivity items, as summarized in Table IV. A comparison of (a) and (b) demonstrates that fine-tuning significantly enhances the domain adaptation ability. Further, the contrast between (b) and (c) (or (d) and (e)) figures out that the prior sensitivity regularization scheme significantly improves SR by $2.4\%$ and PR by $3.2\%$ on LasHeR, highlighting its effectiveness. Similarly, the comparison between (b) and (d) (or (c) and (e)) reveals that rebalancing gradients optimized with diverse transfer data leads to substantial improvements. Notably, the simultaneous application of both prior and transfer sensitivity yields significantly greater improvements than either method alone, proving their complementary nature. While the transfer sensitivity penalty alone yields only modest improvement on FE108, it still contributes to the overall leading performance of our method.
TABLE IV: Ablative study results of the proposed key components. “FFT” refers to full fine-tuning of the backbone; “PS” represents prior sensitivity; “TS” is transfer sensitivity. (a) is the zero-shot performance; (b) serves as our baselines; (e) denotes the complete regularized tuning framework.
Exp. FFT PS TS FE108 DepthTrack LasHeR
@@
Figure 8: Operation-wise eigenvalue spectra for measuring prior sensitivity. Eigenvalues of each operation in OSTrack-B256, obtained during noise probing, are sorted in ascending order and plotted on a logarithmic scale. For each operation, the estimated large eigenvalues show a clear tendency to stabilize, indicating convergent behavior toward the true maximum eigenvalues.
Eigenvalue Budget (Top-$K$) for Prior Sensitivity. In this section, we examine the effect of the eigenvalue count $K$ in Eq. (8). For each operation, we record up to the 100 largest eigenvalues during Rayleigh-quotient probing and compute the prior sensitivity using the top-$K$ of them. As illustrated in Fig. 8, the eigenvalue spectra exhibit a clear saturation behavior in the high-rank regime, indicating convergent behavior toward the true maximum eigenvalues. Consistent with this observation, Table V shows that decreasing $K$ from 20 to 1 has little impact: $K=5$ performs on par with $K=10$ (the best), and even $K=1$ still yields a clear improvement over the baseline. This indicates that the effectiveness of our regularizer depends primarily on the relative ranking of sensitivities rather than their absolute magnitudes. Under this approximation, the key challenge lies in accurately identifying the operations most sensitive to the generalization gap.
@@
Sensitivity Harmony Coefficient $\kappa$. In this section, we evaluate the impact of different sensitivity harmony coefficients, $\kappa$ as defined in Eq. (14), with the detailed results presented in Table VI. By systematically increasing $\kappa$ from 0.1 to 0.9, we observe that a moderately large coefficient (e.g., 0.6) yields the best performance, outperforming both smaller (e.g., 0.3) and larger values (e.g., 0.9). This finding suggests that while both prior sensitivity and transfer sensitivity contribute to performance improvements from different perspectives, a delicate balance between them further optimizes the fine-tuning process, promoting better generalization and adaptability.
@@
Impact of Box Head Tuning. In our default setting, “full fine-tuning” updates the backbone while keeping the box head frozen. To assess the impact of tuning the box head, we also examine a more aggressive variant, complete fine-tuning, where the box head is updated as well. As shown in Table VII, unfreezing the box head further exacerbates overfitting. While our regularization method remains effective, the results reveal an important caveat: unfreezing the box head during cross-modal adaptation can be risky, as it tends to disrupt the modal-agnostic object association knowledge.
@@
Compatibility with PEFT Methods. Existing PEFT methods [70, 17, 51] typically freeze pre-trained parameters and update only a minimal number of additional parameters, which can hinder sufficient optimization. To evaluate the compatibility of our proposed regularization techniques with these methods, we unfreeze their backbones and retrain them with our regularization applied. As shown in Table VIII, our approach significantly enhances ViPT’s performance on VisEvent, yielding gains of $2.1\%$ in SR and $2.0\%$ in PR. Notably, even for the unified tracker UnTrack, our approach achieves superior results on LasHeR: $2.7\%$ in SR and $3.7\%$ in PR. This demonstrates that overly rigid constraints on pre-trained models limit their transfer potential. However, our method negatively impacts SDSTrack, likely because it optimizes the pre-trained parameters, whereas SDSTrack relies on modal-specific adapters trained from scratch.
@@
Compatibility with Single-modality Methods. This work aims to mitigate the misfitting issue when adapting the foundation trackers to downstream tasks. A central question is how the proposed regularization methods perform on single-modality data. To investigate this, we conduct ablation studies highlighting their impact on different modalities. As shown in Table IX, both RGB and auxiliary modalities benefit significantly from our regularization techniques. For example, the RGB and depth modalities achieve F-score gains of $5.3\%$ and $4.0\%$, respectively, on the DepthTrack dataset. Despite substantial distribution differences, our method significantly and consistently enhances the adaptability of auxiliary modalities across various tasks. These findings underscore the importance and necessity of applying constraints when transferring the pre-trained trackers to downstream domains.
@@
Impact of Event Representations. In this work, we focus on constructing suitable frame-like event representations tailored to cross-modal transfer. In this experiment, we keep the same training setup, varying only the input event representations. As shown in Table X, surface-based representations (TSLTD, Time Surface) perform poorly, likely due to the randomness of event timestamps. In contrast, count-based representations (Event Count, Event Frame) offer more robust performance. The time-interpolation-based Event Volume further improves results by utilizing time-weighted and multi-channel methods to preserve spatio-temporal information. Notably, we employ simple color-polarity event images (with ViPT) to align with the RGB pre-trained model, leading to superior performance.
@@
Effectiveness of Smaller Learning Rates. In this paper, we exploit the sensitivity-regularized gradients to bolster the transfer process. A straightforward strategy might involve using a smaller learning rate. To investigate this, we evaluate the effect of reduced learning rates. As shown in Table XI, a smaller learning rate (i.e., $\alpha=10^{-5}$) yields negligible improvement in PR, but at the cost of a decline in SR. Further reduction (i.e., $\alpha=10^{-6}$) leads to overall performance degradation. These findings suggest that merely reducing the learning rate is insufficient to address over-fitting in cross-modality adaptation, while simultaneously limiting transfer potential, as it uniformly suppresses parameter updates without prioritizing those that are sensitive or high-risk.
@@
Exp. $\alpha=10^{-4}$ $\alpha=10^{-5}$ $\alpha=10^{-6}$ Ours
SR 53.2 52.9 52.5 56.3
PR 65.8 66.2 65.4 70.1
@@
TABLE XII: Computational cost of Prior Sensitivity Estimation on OSTrack-B256.
Metric Computational Cost
Operation-wise Model-wise Total Process
FLOPs 104.8 G 15.5 T $\sim$ 39.7 E
Latency $294.5$ ms $43.0$ s 47.8 h ($\sim$2 day)
Memory $164.2$ GB - -
TABLE XIII: Computational complexity and speed analysis on the LasHeR dataset.
@@ IV-D Computational Efficiency
Computational Overhead of Prior Sensitivity. Prior sensitivity is estimated via a group-wise FIM approximation with eigen-decomposition. With the operation-diagonal (group-wise) approximation and Rayleigh-quotient probing, we reduce the computational cost from intractable to a practical level. We further profile this procedure on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, reporting theoretical FLOPs, latency, and peak GPU memory in Table XIV to quantify its efficiency. While this cost is justified by the resulting performance gains, we explicitly acknowledge that—unlike plug-and-play PEFT methods—prior sensitivity estimation introduces additional, non-trivial computation. Importantly, it is a one-shot, offline pre-processing step performed only once before fine-tuning, and thus incurs negligible overhead during subsequent training iterations and inference. Moreover, it can accelerate convergence, effectively offsetting the preprocessing time and potentially reducing the overall time-to-result across the full training lifecycle (see Appendix for details).
Efficiency of Sensitivity-Regularized Fine-tuning. We compare the training efficiency of our method against vanilla fine-tuning. Since prior sensitivity estimation is conducted as a one-time preprocessing step before fine-tuning, its runtime is excluded from the training time. Our method introduces parameter-wise transfer sensitivity quantification with marginal computational overhead, as this sensitivity is derived from the off-the-shelf gradients. While vanilla fine-tuning operates at 37.5 ms per iteration, our method runs at 39.0 ms per iteration, incurring only a modest $4.0\%$ increase in training time.
Inference Cost and Speed. We compare the computational complexity and runtime efficiency of representative baselines, focusing on trackers instantiated with three widely used pre-trained backbones: OSTrack, DropTrack and SUTrack. For a fair comparison, we report both theoretical cost (e.g., parameter count and FLOPs) and practical throughput (FPS) under a unified evaluation protocol. All experiments are conducted on the same hardware platform, equipped with an Intel(R) Xeon(R) 6456C CPU, 256 GB RAM, and a single NVIDIA RTX 3090Ti GPU. As reported in Table XIII, our method supports real-time tracking with 24.8 FPS while delivering top-tier accuracy under the SUTrack setting. When integrated into ViPT (“ViPT+Ours”), it yields both superior speed and performance. Under the DropTrack base model, the resulting tracker still operates at 48.8 FPS, maintaining a respectable speed given its focus on accuracy.
IV-E Limitations and Discussion
From a computational-efficiency perspective, a practical consideration is that prior sensitivity estimation (group-wise FIM approximation with Rayleigh-quotient probing) can still be a non-trivial extra offline procedure compared with plug-and-play PEFT: even after the proposed approximations make it practical, it can still be relatively heavyweight, which can become a hard barrier for large backbones, frequent architecture changes, or rapid ablation cycles. Additionally, the formulation defines the Fisher expectation over the pre-training dataset, so applying the method as-is may require access to (or a governed surrogate for) pre-training data—an assumption that does not always hold under licensing, privacy, or retention constraints—potentially limiting portability across model sources and deployment settings. Addressing these constraints, e.g., by exploring data-free or proxy-based sensitivity estimation—remains an important direction for future work.
@@ V Conclusion
Appendix A Proof of Derivations
A-A Proof of Proposition 1
  1. Bounded FIM Error. Each group approximation $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}$ captures the principal tangent of $\mathcal{F}^{(\theta_{0}^{j})}$ with bounded error. Due to symmetry and orthogonal eigenvectors, the Frobenius norm of the error is derived from the difference in their eigenvalues:
@@ A-A Proof of Proposition 1
  2. Bounded Generalization Gap Error. The approximate FIM $\tilde{\mathcal{F}}^{(\theta_{0})}$ remains close to $\mathcal{F}^{(\theta_{0})}$ as a metric on parameter space. For any parameter difference $\Delta\theta\in\mathbb{R}^{|\theta|}$, let $\varepsilon_{gen}(\mathcal{F}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\mathcal{F}^{(\theta_{0})}\Delta\theta$ denote the generalization gap induced by the true FIM. Likewise $\varepsilon_{gen}(\tilde{\mathcal{F}}^{(\theta_{0})})=\frac{1}{2}\Delta\theta^{T}\tilde{\mathcal{F}}^{(\theta_{0})}\Delta\theta$ is the approximated generalization gap. Then the discrepancy between these distances is bounded in terms of the residual generalization gap error. In particular, one has following formula by definition of the spectral norm:
@@
A-B Proof of Transfer Sensitivity
We state the local assumptions under which the transfer sensitivity bound in Eq. (9)-(12) holds and provide a more explicit derivation.
Assumptions.
Lipschitz-continuous gradients. The loss $L(\theta\mid M)$ is differentiable and $\beta$-smooth in a neighborhood of the current $\theta$, i.e., $\|\nabla L(\theta+\Delta\mid M)-\nabla L(\theta\mid M)\|_{2}\leq\beta\|\Delta\|_{2}$. This implies a first-order Taylor expansion with a quadratic remainder: $L(\theta+\delta\mid M)=L(\theta\mid M)+G^{\top}\delta+r(\delta)$ with $|r(\delta)|\leq\frac{\beta}{2}\|\delta\|_{2}^{2}$.
Bounded single-step gradient magnitude. The per-step gradient magnitude is bounded (a standard assumption in stochastic optimization), which is used when relating gradient sparsity to the amplification of $\|G\|_{2}$ under a fixed $\|G\|_{1}$.
Explicit Perturbation model. $\delta$ and $\delta^{\prime}$ are i.i.d. perturbations around the current update step (as in Eq. (10)).
Setup.
Let $\mathcal{G}=\nabla_{\theta}\mathcal{L}(\theta\mid M)$ and recall
$\epsilon_{\text{ada}}=\mathbb{E}\left[|S\left(\theta,\delta\mid M\right)-S\left(\theta,\delta^{\prime}\mid M\right)\right|],$ (17)
$S(\theta,\delta\mid M)=\mathcal{L}(\theta\mid M)-\mathcal{L}(\theta+\delta\mid M).$ (18)
We model perturbations around a single update step as
$\delta=\alpha(\mathcal{\mathcal{G}}+\xi),\qquad\delta^{\prime}=\alpha(\mathcal{G}+\xi^{\prime}),\qquad$ (19)
where $\xi,\xi^{\prime}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2}I)$, so that $\delta$ and $\delta^{\prime}$ share the same mean step $\alpha\mathcal{G}$ but differ by stochastic perturbations.
Linearization ( Lipschitz continuity).
Assume $L(\theta\mid M)$ is differentiable and locally $\beta$-smooth around $\theta$, i.e., $\|\nabla\mathcal{L}(\theta+\delta\mid M)-\nabla\mathcal{L}(\theta\mid M)\|_{2}\leq\beta\|\delta\|_{2}$. Then a first-order Taylor expansion yields
$\mathcal{L}(\theta+\delta\mid M)=\mathcal{L}(\theta\mid M)+\mathcal{G}^{\top}\delta+r(\delta),\quad|r(\delta)|\leq\frac{\beta}{2}\|\delta\|_{2}^{2},$ (20)
and hence
$S(\theta,\delta\mid M)=-\mathcal{G}^{\top}\delta-r(\delta).$ (21)
Therefore,
$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)=-\mathcal{G}^{\top}(\delta-\delta^{\prime})-(r(\delta)-r(\delta^{\prime})).$ (22)
When perturbations are sufficiently small (so the remainder term is dominated), the leading term is
$S(\theta,\delta\mid M)-S(\theta,\delta^{\prime}\mid M)\approx\mathcal{G}^{\top}(\delta^{\prime}-\delta)=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi).$ (23)
A guaranteed upper bound.
Let $X:=\alpha\,\mathcal{G}^{\top}(\xi^{\prime}-\xi)$. Since $\xi^{\prime}-\xi\sim\mathcal{N}(0,2\sigma^{2}I)$, we have $X\sim\mathcal{N}\!\big(0,\,2\alpha^{2}\sigma^{2}\|\mathcal{G}\|_{2}^{2}\big)$ and thus
$\mathbb{E}|X|\leq\sqrt{\mathbb{E}[X^{2}]}=\sqrt{\mathrm{Var}(X)}=\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2}.$ (24)
Combining Eq. (23) and Eq. (24) gives the following guarantee under the linearized sensitivity:
$\epsilon_{\text{ada}}\approx\mathbb{E}|X|\leq\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2}.$ (25)
This shows that $\epsilon_{\text{ada}}$ scales at most linearly with $\|\mathcal{G}\|_{2}$.
Tighter closed form.
Moreover, for a zero-mean Gaussian $X\sim\mathcal{N}(0,s^{2})$, $\mathbb{E}|X|=s\sqrt{2/\pi}$. Hence,
$\mathbb{E}|X|=\sqrt{\frac{2}{\pi}}\sqrt{2\alpha^{2}\sigma^{2}\|\mathcal{G}\|_{2}^{2}}=\frac{2\alpha\sigma}{\sqrt{\pi}}\,\|\mathcal{G}\|_{2}\leq\alpha\sqrt{2\sigma^{2}}\,\|\mathcal{G}\|_{2},$ (26)
where the last inequality holds since $2/\sqrt{\pi}\leq\sqrt{2}$.
Deviation from linearization.
From Eq. (22)–(20), the approximation error is controlled by the Taylor remainders:
$\Big|\epsilon_{\text{ada}}-\mathbb{E}|X|\Big|\leq\mathbb{E}|r(\delta)-r(\delta^{\prime})|\leq\frac{\beta}{2}\mathbb{E}\big(\|\delta\|_{2}^{2}+\|\delta^{\prime}\|_{2}^{2}\big)=\beta\,\mathbb{E}\|\delta\|_{2}^{2},$ (27)
so the linear scaling is tightest in locally smooth regions and for sufficiently small perturbations.
Appendix B Additional ablation studies
B-A Computational Overhead of Prior Sensitivity under Larger Backbones
In addition to the pre-trained trackers used in this work (e.g., OSTrack-B256 with a ViT-B backbone), we further investigate the FLOPs and memory costs of the eigen-decomposition-based prior sensitivity estimation under larger backbones, such as ViT-L and Swin-L. Based on empirical measurements conducted on a cluster of 8 NVIDIA RTX 3090 Ti GPUs, the specific computational overhead introduced by FIM estimation for ViT-L and Swin-L is detailed in Tab. XIV. It is important to note that the total cost of the offline FIM process is data-dependent, scaling with the size of the pre-training dataset and the number of iterations required for convergence given the model capacity. Consequently, rather than providing a variable total estimate, we report the deterministic per-group and per-iteration costs to accurately reflect the computational costs.
TABLE XIV: Computational cost of FIM eigen-decomposition on ViT-L and Swin-L ($\times256$ input resolution).
Model Metric Granularity Computational Cost
Per-Group (Operation-wise) Per-Iteration (Model-wise)
ViT-L FLOPs Theoretical Ops 170.2 GFLOPs 49.4 TFLOPs
Latency Wall-clock Time 452.5 ms 131.2 s
[…]
Swin-L FLOPs Theoretical Ops 323.1 GFLOPs 93.7 TFLOPs
Latency Wall-clock Time 687.8 ms 206.3 s
[…]
B-B Training Efficiency of SRFT
We explicitly acknowledge that unlike plug-and-play PEFT methods (e.g., Prompt Tuning, Adapters, LoRA), which incur zero preprocessing cost via random initialization, our SRFT involves an offline FIM estimation phase. However, we argue that this offline computation represents a strategic “pay once, benefit long-term” investment. This upfront cost is justified by the following characteristics over PEFT:
1. Accelerated Convergence: The FIM estimation is a one-shot process performed on the pre-training datasets. The resulting sensitivity map is then frozen and reused for all subsequent multi-modal tracking tasks. Furthermore, the pre-trained weight acts as a “smart initialization” compared to the random initialization of PEFT. By guiding the optimization along the most reliable directions from the start, SRFT significantly accelerates convergence. As shown in Tab. XV and Fig. 11, SRFT requires fewer optimal iterations and epochs to reach optimal performance compared to PEFT-based methods, effectively offsetting the preprocessing time and potentially achieving a lower total time-to-result for the full training lifecycle.
2. Zero Inference Latency: This is of paramount importance for real-time applications. As detailed in the original manuscript, the proposed regularization techniques are applied exclusively during training. Consequently, they impose no storage overhead and zero additional inference latency during testing, which is crucial for real-time tracking scenarios.
3. Superior Performance: Finally, the computational cost translates directly into performance gains. By leveraging the global sensitivity information, SRFT consistently outperforms PEFT methods in accuracy (as detailed in the main paper tables), proving that the offline overhead yields a high return on performance.
TABLE XV: Comparison of preprocess time, trainable parameter and training time across different multi-modal tracking methods. All experiments are conducted based on the pre-trained OSTrack-B256. “Preprocessing” refers to prior sensitivity estimation. The training epochs (listed sequentially from left to right) correspond to the FE108, VisEvent, CoeSot, LasHeR, and DepthTrack datasets, respectively. Note that UnTrack employs a joint training strategy on the VisEvent, LasHeR, and DepthTrack datasets.
Method Trainable Parameter (M) Preprocess Time (h) Total Epochs Training Time (h) Total Time (h)
ViPT (Prompt) 0.8 - 300 $(60\times 5)$ $15.3min\times 300\approx 76.5\text{ h}$ 76.5
ViPT + SRFT 86.4 47.8 h 100 $(20\times 5)$ $15.9min\times 100\approx 26.5\text{ h}$ 74.3
SDSTrack (Adapter) 14.8 - 205 $(50\times 3+40+15)$ $21.9min\times 205\approx 74.8\text{ h}$ 74.8
SDSTrack + SRFT 100.6 47.8 h 85 $(20\times 3+15+10)$ $22.8min\times 85\approx 32.3\text{ h}$ 80.1
UnTrack (LoRA) 6.6 - 80 $18.5min\times 80\approx 19.3\text{h}$ 19.3
SDSTrack + SRFT 92.4 47.8 h $40$ $19.4min\times 40\approx 10.3\text{ h}$ 58.1
Figure 11: Comparison of training loss for different multi-modal tracking methods on DepthTrack dataset. Note that all methods share the same loss function and adopt pre-trained OSTrack initialization. Evidently, our method converges in fewer training iterations and reaches a lower loss value, effectively mitigating underfitting. Moreover, the training dynamics are noticeably more stable, exhibiting reduced oscillation and improved consistency throughout fine-tuning.
Figure 12: Schematic illustration of different sensitivity fusion scheduling strategies. We compare the evolution of the prior sensitivity fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) using Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial sensitivity harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
B-C Different Sensitivity Fusion Scheduling Strategies
Our design is motivated by the stability–plasticity trade-off in cross-domain fine-tuning: effective transfer requires balancing stability (preserving critical pre-trained knowledge) and plasticity (adapting to the target domain). In the paper, we explicitly define a dynamic linear schedule to harmonize the two sensitivities: at the beginning, the prior sensitivity has weight $\kappa$, and as training proceeds the transfer sensitivity gradually increases until they reach the same weight $\kappa$ at the end. This is formalized by Eq. (14), where the combined sensitivity is a convex combination with linear dependence on training progress $\frac{t}{T}$.
As summarized in the Fig. 12 and Tab. XVI, we additionally compare exponential, cosine, piecewise (4-stage), and linear (ours) schedules on FE108 under sensitivity harmony coefficient $\kappa\in\{0.3,0.6,0.9\}$. To ensure a fair comparison of the trajectory effect induced by different schedules, we enforce the same initial and final weights for all schedules, so that the only difference lies in how the weights evolve over training. Overall, the linear schedule is consistently best or tied-best. When $\kappa=0.6$ (our default), both the weight evolution and the final performance are very close across different schedules, indicating the schedule choice is not overly sensitive in this moderate regime. However, when $\kappa$ becomes too small or too large, the alternative schedules tend to be more fragile. Mechanistically, exponential/cosine may over-emphasize one sensitivity at certain phases (e.g., too aggressive early transfer or overly conservative late adaptation), whereas piecewise scheduling introduces abrupt weight jumps that can destabilize optimization. In contrast, the linear schedule provides a monotonic and smooth transition with a constant rate of change, which empirically leads to more stable and robust behavior when $\kappa$ deviates from the moderate setting.
TABLE XVI: Comparison of different sensitivity fusion schedules on FE108 dataset. We compare the prior sensitivity fusion weight $w_{p}$ over the training process ($\frac{t}{T}$) with Linear (Ours), Cosine, Exponential, and Piecewise functions across varying initial sensitivity harmony coefficient settings ($\kappa\in\{0.3,0.6,0.9\}$).
Schedule Prior Sensitivity Weight Transfer Sensitivity Weight $\kappa$ SR PR
Exponential $w^{p}=\kappa\left(\frac{1-\kappa}{\kappa}\right)^{\frac{t}{T}}$ $w^{t}=1-w^{p}$ 0.3 65.7 94.2
0.6 67.2 96.3
0.9 66.4 94.1
Cosine $w^{p}=(1-\kappa)+\frac{2\kappa-1}{2}\left(1+\cos\left(\frac{t\pi}{T}\right)\right)$ $w^{t}=1-w^{p}$ 0.3 66.6 95.0
0.6 67.3 96.4
0.9 66.7 95.2
Piecewise (4-Stage) $w^{p}=\kappa+\frac{1-2\kappa}{3}\lfloor\frac{4t}{T}\rfloor$ $w^{t}=1-w^{p}$ 0.3 66.8 94.9
0.6 67.1 96.0
0.9 66.5 95.3
Linear (Ours) $w^{p}=\kappa+(1-2\kappa)\frac{t}{T}$ $w^{t}=1-w^{p}$ 0.3 66.9 95.1
0.6 67.4 96.5
0.9 66.9 95.5
B-D Visual Comparison